OpenStudy (anonymous):

the equation x/-4+3=-4

OpenStudy (bahrom7893):

x - 12 = 16 x = 28

OpenStudy (saifoo.khan):

28

OpenStudy (anonymous):

x/-4 + 3 = -4 - 3 - 3 x/-4 = -7 * -4 * -4 x = 28